(1 point) If E[X] = -1 and Var(X) = 4, then E[(2 + 4X)?] = and Var(2 + 3X) =
Answer #1

E(2 + 4X)2 E ( 4 + 2 * 2 * 4 X + 16 X2)

= 4 + 16 E(X) + 16 E (X2)

= 4 + 16 E(X) + 16 [ Var(X) + E(X)2 ] , [ Since Var(X) = E (X2) - E(X)2 ]  

= 4 + 16 (-1 ) + 16 [ 4 + 1]

= 68

Var ( 2 + 3 X) = Var(2) + 32 Var (X)  

= 0 + 9 * 4

= 36